OpenStudy (anonymous):

.84 repeating as a fraction

OpenStudy (anonymous):

Is that the question?

OpenStudy (anonymous):

close to 59202/70000

OpenStudy (anonymous):

but i don't know the exact fraction

OpenStudy (anonymous):

38/45 is even better

OpenStudy (amistre64):

continued fraction maybe? .84 = 84/100 \[\frac{84}{100} =\frac{21}{25}=\cfrac{1}{\left(\cfrac{25}{21}\right)} =\cfrac{1}{1+\frac{4}{21}}\] \[\cfrac{1}{1+\frac{4}{21}}=\cfrac{1}{1+\cfrac{1}{\left(\frac{21}{4}\right)}}=\cfrac{1}{1+\cfrac{1}{5+\cfrac{1}{4}}}\]

OpenStudy (amistre64):

oh .8484848484...

OpenStudy (amistre64):

84/99

OpenStudy (amistre64):

let A = .848484... 100A = 84.848484... 100A - A = 99A = 84